# When Text Embedding Meets Large Language Model: A Comprehensive Survey

### v2

@@ 7.1.1 Cross-Lingual Text Embedding
Some LLMs are pre-trained on large multi-lingual corpora, which leads us to expect higher-quality cross-language text embeddings using LLMs. Some exploratory work has attempted to introduce LLM into cross-language scenarios and has shown remarkable potential. E5-Mistral [153] demonstrates the great potential of LLM for synthesizing data and being a backbone, while UDEVER [190] shows the powerful generalization capabilities of multilingual LLM, even when fine-tuned using only English text.
In addition, both OpenAI and Cohere offer multilingual embedding services; Cohere blogged about the second phase of its model training using additional supervised signals generated by LLM  [^19] https://cohere.com/blog/introducing-embed-v3; and OpenAI’s currently disclosed tech line [189] and up to 3,072 embedding dimensions give us a reasonable suspicion that the LLM backs the service as an encoder.
@@ 7.1.2 Cross-Modal Text Embedding
In the cross-modal domain, joint language-vision embedding learning [309, 310, 311] has received the most attention. Its applications have gradually expanded from image-text retrieval [312, 313] to multiple tasks such as composed image retrieval [314, 315], multi-modal document retrieval [316] and multi-modal knowledge retrieval [317, 318]. VISTA [319] generates high-quality instructions and captions using LLMs and achieves substantial improvements on various cross-modal retrieval tasks. For example, E5-V [320] fine-tuning MLLMs with text-modal contrastive learning only, firstly showing the potential of MLLMs as a multi-modal unified encoder; GME [321] synthesized large-scale and high-quality data for fine-tuning MLLM through the pipeline, showing strong performance on uni-modal (even visual modal) and multi-modal by training the model on text pairs alone. LLaVE [322]
@@ 7.2 Task-Specific Text Embedding
Therefore, we can convert the hope from universal text embeddings to universal text encoders $F$, which can be expressed as
@@ 7.2.1 Instruction-Following Embedding
In instruction-following embedding learning, $F$ is divided into two parts: $F=f\circ p_{t}$, where $p_{t}$ is a prompt template related to task $t$. In the beginning, prompts are used to align the fine-tuning targets with the pre-trained targets of the PLMs [327]. Then, similar instruction tuning [328] is proposed to adjust LLMs to be more adaptable to the format of task-oriented conversations. Instruction-following embedding learning follows the idea of instruction tuning but models all tasks into a contrastive paradigm. TART [51] firstly validate the feasibility of this idea to information retrieval, while Instructor [67] fine-tuning GTR on 330 NLP tasks and find that including instructions would alleviate conflicts between tasks compared to traditional multi-task learning. In the zero-shot evaluation, the model fine-tuned with instructions shows good generalization to new instructions and datasets. The findings related to this technique are still iterating: [164] add instruction only on evaluation and demonstrate its effectiveness; [191] demonstrate that question-and-answer pairs can significantly improve a model’s ability to comply with instructions under a particular training method.

### v3

@@ 7.1.1 Cross-Lingual Text Embedding
LLM-Based Embedder
The common backbones used by LLM-based embedders, such as Mistral and Qwen2, are pre-trained on large multi-lingual corpora, which leads us to expect higher-quality cross-language text embeddings using LLMs. UDEVER [196] shows the powerful embedding generalization capabilities of multi-lingual LLM, even when fine-tuned using only English text. Subsequent embedders based on Mistral or Qwen, such as GTE-Qwen2, GirtLM, and E5-Mistral, have achieved extraordinary performance on MTEB after contrastive learning on large-scale multi-lingual corpora. Specifically, E5-Mistral [157] synthesized a large-scale training corpus using LLM to enhance data diversity; GTE-Qwen2 [198] introduced weakly-supervised-supervised two-stage contrastive learning, which fully utilized the massive noisy web corpus. xVLM2Vec [311] extends cross-modal embeddings from MLLM-based embedders further in cross-language settings through self-knowledge distillation. Gemini-Embedding [312] integrates the experience of synthetic data and two-stage contrastive learning and, based on this, utilizes the technique of model merging [313] to further improve the performance of embedding on multi-lingual tasks.
Commercial Service
Both OpenAI and Cohere offer multilingual embedding services; Cohere blogged about the second phase of its model training using additional supervised signals generated by LLM  [^19] https://cohere.com/blog/introducing-embed-v3; and OpenAI’s currently disclosed tech line [195] and up to 3,072 embedding dimensions give us a reasonable suspicion that the LLM backs the service as an encoder.
Evaluation
Cross-lingual embedding as a long-standing research area has diverse downstream task evaluation benchmarks, such as MASSIVE [314] for Classification, STS-17 [114] and STS-22 [121] for semantic text similarity, MKQA [315] for question answering, Mr.TyDi [58], XOR-Retrieve [316], XTREME-UP [317] and MIRACL [57] for text retrieval, CodeSearchNet [318] for code search. Newly proposed evaluation benchmarks have recently become more difficult or compounded, such as MLDR [56] for long document retrieval, mFollowIR [319] for instruction following; MTEB(Code) [^20] http://mteb-leaderboard.hf.space/?benchmark_name=MTEB%28Code%2C+v1%29 and MMTEB [320] for universal multi-lingual text embedding.
Table 6: The overview of MLLM-based embedders. The paradigm part is shortened as follows: multi-modal supervised contrastive learning (MSCL), text supervised contrastive learning (TSCL), masked next token prediction (MNTP), and next token prediction (NTP).
Method Model Architecture Training Evaluation
LLM (Encoder) Pooling Attention Projector PEFT Input Format Paradigm Task
E5-V [321] LLaVa-1.6 Last Causal $\times$ LoRA Prompt MSCL Flickr30K, COCO, CIRR, FashionIQ, STS
LLM2CLIP [322] LLaMA+ViT Mean Bi $\surd$ LoRA Prompt MNTP→MSCL Flickr30K, COCO, ShareGPT4V, Urban-1k, DOCCI
LamRA-Ret [323] Qwen2-VL Special Last Causal $\times$ LoRA Instruction+Prompt TSCL->MSCL M-BEIR
DSE [324] Phi-3-V Special Last Casual $\times$ LoRA Prompt MSCL NQ, SlideVQA-open
InstructCIR [325] LLaVA-Phi-3-mini Special Last Causal $\times$ LoRA Instruction MSCL CIRCO, CIRR, FashionIQ
VladVA [326] LLaVa-1.5 Special Last Causal $\times$ LoRA Prompt MSCL+NTP Flickr30K, COCO, SugarCrepe, SugarCrepe++
MMRet-MLLM [327] LLaVa-1.6 Special Last Causal $\times$ LoRA Instruction MSCL CIR, MMEB
GME [328] Qwen2-VL Special Last Causal $\times$ LoRA Instruction MSCL UMRB, BEIR, M-BEIR, ViDoRe
VLM2Vec [329] Phi-3.5-V, LLaVa-1.6 Last Causal $\times$ LoRA / $\times$ Instruction MSCL MMEB
UniSE-MLLM [330] Qwen2-VL Special Last Causal $\times$ LoRA Instruction MSCL MVRB
MM-Embed [331] LLaVa-1.6+NV-Embed Post Bi $\times$ LoRA Instruction MSCL->TSCL M-BEIR + MTEB
ColPali [332] PaliGemma - (Late Interaction) Prefix $\surd$ LoRA Instruction MSCL ViDoRe
LLaVE [333] LLaVA-OV, AquilaVL Last Causal $\times$ $\times$ Instruction MSCL MMEB
@@ 7.1.2 Cross-Modal Text Embedding
In the cross-modal domain, joint language-vision embedding learning [334, 335, 336] has received the most attention. Its applications have gradually expanded from image-text retrieval [337, 338] to multiple tasks such as composed image retrieval [339, 340], multi-modal document retrieval [341] and multi-modal knowledge retrieval [342, 343], etc. Before the advent of multimodal large language models (MLLMs), ONE-PEACE [344] raised the number of parameters to 4B based on customizing the architecture, enabling a unified representation of images, text, and audio.
MLLM-Augmented Multi-Modal Embedding
Unlike text data, multimodal data is smaller-scale and more noisy, hindering the development of multi-modal Embedding. Recent studies have shown that using LLM to augment multi-modal data while employing good filtering will effectively improve the performance of multi-modal embedding across multiple tasks. For example, VISTA [345] generates high-quality instructions and captions with the powerful OpenAI LLM and Stable Diffusion, achieving substantial improvements on various cross-modal retrieval tasks. SPN [346] uses MLLMs and the pre-trained image encoder to synthesize both positive and negative examples for the composed image retrieval task, successfully improving the performance of multiple baseline models. To train the MLLM-based embedder, the larger scale instruction multi-modal dataset, such as MegaPairs [327] and VIRA [330], are synthesized with LLM-MLLM pipelines. MLLM-Embedder trained on these datasets show strong performance on uni-modal (even visual modal) and multi-modal tasks.
MLLM-Based Embedder
As shown in Table 6, many attempts at MLLM-based embedders have been inspired by the successful experience of LLM-based embedders. For example, E5-V [321] follow the prompt style of PromptEOL [170] and fine-tune MLLMs with text contrastive learning only, firstly showing the potential of MLLMs as a multi-modal unified embedder. LLM2CLIP [322] trains a LLM-based embedder following LLM2Vec [168] and replaces the text encoder of CLIP with this LLM-based embedder, improving the performance of the CLIP model on multiple tasks effectively. MM-Embed [331] introduces the incremental text contrastive learning stage after traditional multi-modal contrastive learning, showing strong performance in both text and multi-modal embedding tasks. VladVA [326] uses next token prediction to increase the task difficulty for long text, preventing early gradient dissipation in the optimization process. In addition, many effective strategies in text embedding methods are followed, such as instruction tuning [323, 329], hard negative mining [328, 331], scaling negative number [333], etc. Except for MLLM-based embedders for universal multi-modal and uni-modal tasks, some MLLM-based embedders focus on the optimization of complex tasks in real scenarios. For example, DSE [324], UniSE-MLLM [330], and ColPali [332] focus on tasks related to document screenshot understanding. They leverage MLLM’s powerful comprehension capabilities to simplify the original complex embedding process while improving performance and efficiency. InstructCIR [325] focuses on composed image retrieval, which achieves better instruction-following capability by two-stage contrastive learning while progressive freezing the modules in the MLLM.
Evaluation
As the number of MLLM-based embedders increased, evaluation benchmarks with different focuses were proposed, such as M-BEIR [336] for multi-modal retrieval, MVRB [330] for multi-modal document understanding, MMTB [329] for multi-modal universal embedding and UMRB [328] for both uni-modal and multi-modal universal embedding. However, we find that many works do not indicate whether their evaluation is a true zero-shot out-of-domain or not, and there was a risk of unfair performance comparison. This is the reason why we do not list “evaluation-setting” in Table 6 as we do in Table 4. We call on subsequent work to filter the training data strictly to prevent overlapping training-evaluation data, providing a fair performance benchmark for the community.
@@ 7.2 Task-Specific Text Embedding
Therefore, we can convert the hope from universal text embeddings to universal text encoders $F$, which is expressed as
@@ 7.2.1 Instruction-Following Embedding
In instruction-following embedding learning, $F$ is divided into two parts: $F=f\circ p_{t}$, where $p_{t}$ is a prompt template related to task $t$. In the beginning, prompts are used to align the fine-tuning targets with the pre-trained targets of the PLMs [351]. Then, similar instruction tuning [352] is proposed to adjust LLMs to be more adaptable to the format of task-oriented conversations. Instruction-following embedding learning follows the idea of instruction tuning but models all tasks into a contrastive paradigm. TART [51] firstly validates the feasibility of this idea to information retrieval, while Instructor [67] fine-tuning GTR on 330 NLP tasks and found that including instructions would alleviate conflicts between tasks compared to traditional multi-task learning. In the zero-shot evaluation, the model fine-tuned with instructions shows good generalization to new instructions and datasets. The findings related to this technique are still iterating: LLM2Vec [168] add instruction only on evaluation and demonstrate its effectiveness; Inbedder [197] demonstrate that question-and-answer pairs can significantly improve a model’s ability to comply with instructions under a particular training method.
@@ 8 Conclusion
Contribution
Zhijie Nie was responsible for the outline and versioning of this survey, writing Section 1, 2, 6, and 7, and involved in writing Section 4; Zhangchi Feng was responsible for formatting the dissertation of this survey, collecting the dissertation and writing for Section 5.1; Mingxin Li was responsible for collecting the dissertation and writing for Section 3; Cunwang Zhang was responsible for collecting the dissertation and writing for Section 5.2; Yanzhao Zhang was involved in writing Sections 2 and 4; and Dingkun Long was involved in writing Sections 4; Rizhong Zhang provided financial support and coordinated the review.
We are grateful for the efforts made by Hailang Huang for the advance preparation of this work. We thank Raghuveer Thirukovalluru for his suggestions for improving our paper.